Loookout station A is 15 km west of station B. The bearing from A to a fire directly south of B is S 37 degrees 50' E. How far is the fire from B?
Answer by stanbon(75887)

Loookout station A is 15 km west of station B. The bearing from A to a fire directly south of B is S 37 degrees 50' E. How far is the fire from B?
-----------
Draw the picture. You have a right triangle at pt. B.
One leg of that angle is 15 km.
The other leg is the distance you are looking for (call it "x")
and it is opposite the given angle.
EQUATION:
tangent (37 deg. 50')= x/15
x=15*tan(37.83 deg)=15(0.7765...)=11.65 km
Cheers,
Stan H.